# Arbitrage-free catastrophe reinsurance valuation for compound dynamic contagion claims

### v1

@@ 3.1 Equivalent martingale measures
We now describe how the probability laws of $N_{t}$ and $C_{t}$ change after changing probability measures by the Esscher transform. Firstly, from Theorem 3.1, $\{\mathrm{e}^{K(t)}\theta^{N_{t}}\mathrm{e}^{B(t)\lambda_{t}}~{}\mathrm{e}^{-\nu C_{t}}{\mathrm{e}^{\phi\Lambda_{t}}\psi^{M_{t}}}\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P})$-martingale. Then, for all $t\in{\cal T}$,
@@
Consider the following (normalized) $\mathbb{F}$-adapted process $\{\zeta_{t}\}_{t\in{\cal T}}$:
@@
It is clear that by its definition in Eq. 3.10 and Theorem 3.1, $\{\zeta_{t}\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P})$-martingale. For each $t\in{\cal T}$, $\zeta_{t}>0$, $\mathbb{P}$-a.s. Furthermore, $\mathbb{E}_{0}[\zeta_{t}]=\zeta_{0}=1$. Consequently, a new probability measure $\mathbb{P}^{\ast}$, which is equivalent to the original probability measure $\mathbb{P}$ on ${\cal F}_{T}$, can be defined by putting:
$\frac{d\mathbb{P}^{\ast}}{d\mathbb{P}}\bigg{|}_{{\cal F}_{T}}:=\zeta_{T}.$
Let $\tilde{\mathcal{A}}$ denote the infinitesimal generator of the joint process $\{\left(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t\right)\}_{t\in{\cal T}}$ under the new probability measure $\mathbb{P}^{\ast}$. The key to study how the probability laws of $N_{t}$ and $C_{t}$ change after changing probability measures by the Esscher transform is to obtain the infinitesimal generator $\tilde{\mathcal{A}}$. This is to be achieved in Theorem 3.3 and Corollary 3.1 to be presented in the sequel.
@@
Let ${\cal D}({\cal A})$ and ${\cal D}(\tilde{\mathcal{A}})$ denote the domains of the infinitesimal generators ${\cal A}$ and $\tilde{\mathcal{A}}$, respectively. Then, for any function $\tilde{f}\in{\cal D}({\cal A})\cap{\cal D}(\tilde{\mathcal{A}})$,
$\tilde{\mathcal{A}}\tilde{f}\left(\lambda,n,c,{m},{\Lambda},t\right)$
@@
$\quad+\theta\hat{j}(\nu)\hat{g}(-B(t))~{}\lambda\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right]$
$\quad+\tilde{\rho}(t)\left[\int_{0}^{\infty}\tilde{f}\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right].$ (3.11)
@@
$\mathrm{e}^{K(t)}\theta^{n}\mathrm{e}^{B(t)\lambda}~{}\mathrm{e}^{-\nu c}{\mathrm{e}^{\phi\Lambda}\psi^{m}}\tilde{f}\left(\lambda,\Lambda,n,c,{m},{\Lambda},t\right),$
into the generator in Section 2.2, we have
$\tilde{\mathcal{A}}\tilde{f}\left(\lambda,n,c,{m},{\Lambda},t\right)$
$=\frac{\mathcal{A}\bigg{\{}\mathrm{e}^{K(t)}\theta^{n}\mathrm{e}^{B(t)\lambda}~{}\mathrm{e}^{-\nu c}{\mathrm{e}^{\phi\Lambda}\psi^{m}}\tilde{f}\left(\lambda,\Lambda,n,c,{m},{\Lambda},t\right)\bigg{\}}}{\mathrm{e}^{K(t)}\theta^{n}\mathrm{e}^{B(t)\lambda}~{}\mathrm{e}^{-\nu c}{\mathrm{e}^{\phi\Lambda}\psi^{m}}}$
$=\frac{\partial\tilde{f}}{\partial t}+\bigg{(}K^{\prime}(t)+B^{\prime}(t)\lambda\bigg{)}\tilde{f}+\delta\left(a-\lambda\right)\left(\frac{\partial\tilde{f}}{\partial\lambda}+B(t)\tilde{f}\right)+{\lambda\left(\frac{\partial\tilde{f}}{\partial\Lambda}+\phi\tilde{f}\right)}$
$\quad+\lambda\left[\theta\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathrm{e}^{B(t)y}\mathrm{e}^{-\nu\xi}\mathop{}\!\mathrm{d}G(y)\mathop{}\!\mathrm{d}J(\xi)-\tilde{f}\right]$
$\quad+\rho\left[\psi\int_{0}^{\infty}\tilde{f}\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathrm{e}^{B(t)x}\mathop{}\!\mathrm{d}H(x)-\tilde{f}\right]$
$=\frac{\partial\tilde{f}}{\partial t}+\bigg{(}K^{\prime}(t)+B^{\prime}(t)\lambda\bigg{)}\tilde{f}+\delta\left(a-\lambda\right)\left(\frac{\partial\tilde{f}}{\partial\lambda}+B(t)\tilde{f}\right)+{\lambda\left(\frac{\partial\tilde{f}}{\partial\Lambda}+\phi\tilde{f}\right)}$
$\quad+\lambda\left[\theta\hat{j}(\nu)\hat{g}(-B(t))\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\frac{\mathrm{e}^{B(t)y}}{\hat{g}(-B(t))}\mathop{}\!\mathrm{d}G(y)\frac{\mathrm{e}^{-\nu\xi}}{\hat{j}(\nu)}\mathop{}\!\mathrm{d}J(\xi)-\tilde{f}\right]$
$\quad+\rho\left[\psi\hat{h}(-B(t))\int_{0}^{\infty}\tilde{f}\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\frac{\mathrm{e}^{B(t)x}}{\hat{h}(-B(t))}\mathop{}\!\mathrm{d}H(x)-\tilde{f}\right].$
@@
$\tilde{\mathcal{A}}\tilde{f}\left(\lambda,n,c,{m},{\Lambda},t\right)$
$=\frac{\partial\tilde{f}}{\partial t}+\bigg{(}K^{\prime}(t)+B^{\prime}(t)\lambda\bigg{)}\tilde{f}+\delta\left(a-\lambda\right)\left(\frac{\partial\tilde{f}}{\partial\lambda}+B(t)\tilde{f}\right)+{\lambda\left(\frac{\partial\tilde{f}}{\partial\Lambda}+\phi\tilde{f}\right)}$
$\quad+\lambda\left[\theta\hat{j}(\nu)\hat{g}(-B(t))\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right]$
$\quad+\rho\left[\psi\hat{h}(-B(t))\int_{0}^{\infty}\tilde{f}\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right]$
@@
$\quad+\lambda\left[\theta\hat{j}(\nu)\hat{g}(-B(t))\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right]$
$\quad+\rho\left[\psi\hat{h}(-B(t))\int_{0}^{\infty}\tilde{f}\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right]$
$=\bigg{[}\rho\left[1-{\psi}\hat{f}(-B(t))\right]+\bigg{(}1-\theta~{}{\hat{j}(\nu)}~{}\hat{g}(-B(t))\bigg{)}\lambda\bigg{]}\tilde{f}+\frac{\partial\tilde{f}}{\partial t}+\lambda\frac{\partial\tilde{f}}{\partial\Lambda}+\delta\left(a-\lambda\right)\frac{\partial\tilde{f}}{\partial\lambda}$
$\quad+\lambda\left[\theta\hat{j}(\nu)\hat{g}(-B(t))\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right]$
$\quad+\rho\left[\psi\hat{h}(-B(t))\int_{0}^{\infty}\tilde{f}\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right]$
@@
$\quad+\lambda\left[\theta\hat{j}(\nu)\hat{g}(-B(t))\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\theta~{}{\hat{j}(\nu)}~{}\hat{g}(-B(t))\tilde{f}\right]$
$\quad+\rho\left[\psi\hat{h}(-B(t))\int_{0}^{\infty}\tilde{f}\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-{\psi}\hat{f}(-B(t))\tilde{f}\right]$
$\quad+\theta\hat{j}(\nu)\hat{g}(-B(t))\lambda\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right]$
$\quad+\psi\hat{h}(-B(t))~{}\rho\left[\int_{0}^{\infty}\tilde{f}\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right].$
@@
    The self-exciting intensity function $\lambda_{t}$ has changed to $\theta\hat{j}(\nu)\hat{g}(-B(t))~{}\lambda_{t}$, which depends on time;
@@
$\tilde{\mathcal{A}}\tilde{f}\left(\tilde{\lambda},n,c,{m},{\tilde{\Lambda}},t\right)$
@@
$\quad+\tilde{\lambda}\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\tilde{\lambda}+u,n+1,c+\xi,{m},{\tilde{\Lambda}},t\right)\mathop{}\!\mathrm{d}\tilde{G}\left(\frac{u}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t\right)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right]$
$\quad+\tilde{\rho}(t)\left[\int_{0}^{\infty}\tilde{f}\left(\tilde{\lambda}+v,n,c,{m+1},{\tilde{\Lambda}},t\right)\mathop{}\!\mathrm{d}\tilde{H}\left(\frac{v}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t\right)-\tilde{f}\right].$ (3.12)
@@
$\tilde{\mathcal{A}}\tilde{f}\left(\tilde{\lambda},n,c,{m},{\tilde{\Lambda}},t\right)$
@@
$\quad+\tilde{\lambda}\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\tilde{\lambda}+\theta\hat{j}(\nu)\hat{g}(-B(t))y,n+1,c+\xi,{m},{\tilde{\Lambda}},t\right)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right]$
$\quad+\tilde{\rho}(t)\left[\int_{0}^{\infty}\tilde{f}\left(\tilde{\lambda}+\theta\hat{j}(\nu)\hat{g}(-B(t))x,n,c,{m+1},{\tilde{\Lambda}},t\right)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right].$
@@
$a$ $\rightarrow\ \theta\hat{j}(\nu)\hat{g}(-{B(t)})~{}a;$
$\rho$ $\rightarrow\ \psi\hat{h}(-{B(t)})~{}\rho;$
$h(v)$ $\rightarrow\ \frac{\tilde{h}\left(\frac{v}{\theta\hat{j}(\nu)\hat{g}(-{B(t)})};t\right)}{\theta\hat{j}(\nu)\hat{g}(-{B(t)})};$
$g(u)$ $\rightarrow\ \frac{\tilde{g}\left(\frac{u}{\theta\hat{j}(\nu)\hat{g}(-{B(t)})};t\right)}{\theta\hat{j}(\nu)\hat{g}(-{B(t)})};$
$j(\xi)$ $\rightarrow\ \tilde{j}(\xi),$
@@
$\tilde{\mathcal{A}}\tilde{f}\left(\tilde{\lambda},n,c,{m},{\tilde{\Lambda}},t\right)$
@@
$\quad+\tilde{\lambda}\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}\left(\tilde{\lambda}+u,n+1,c+\xi,{m},{\tilde{\Lambda}},t\right)\mathop{}\!\mathrm{d}\tilde{J}(\xi)~{}\frac{\tilde{g}\left(\frac{u}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t\right)}{\theta\hat{j}(\nu)\hat{g}(-B(t))}\mathop{}\!\mathrm{d}u-\tilde{f}\right]$
$\quad+\tilde{\rho}(t)\left[\int_{0}^{\infty}\tilde{f}\left(\tilde{\lambda}+v,n,c,{m+1},{\tilde{\Lambda}},t\right)\frac{\tilde{h}\left(\frac{v}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t\right)}{\theta\hat{j}(\nu)\hat{g}(-B(t))}\mathop{}\!\mathrm{d}v-\tilde{f}\right].$
@@
Note that the parameters of a CDCP under the original measure $\mathbb{P}$ are all constants, whereas the parameters of a CDCP under the new measure $\mathbb{P}^{\ast}$ become time-varying (except for the intensity-decay rate $\delta$), i.e., a time-inhomogeneous CDCP.
@@
In a hardening reinsurance market, it may not be unreasonable for a reinsurer to calculate the premium of a catastrophe reinsurance contract in a more conservative and prudent way. To illustrate this situation in our current modelling set up, certain assumptions are imposed on $\theta,{\psi},\nu,B(t)$ such that there would be more external-exciting jump arrivals in a given period of time, a higher value of self-exciting intensity, higher values of their jump sizes, as well as a larger size of claim or loss. To take on these additional risks in the hardening market, a reinsurer requires a higher compensation. Furthermore, it may not be unreasonable to postulate that the reinsurer aims to maximize their shareholders’ wealth via earning profits. To this end, the reinsurer may charge a higher premium than the “break-even” premium. The latter is given by the expected claims evaluated under the original (real-world) probability measure. In the hardening market, $\theta,{\psi},\nu,B(t)$ may be thought of as security loading factors to be evaluated by which positive gross premium should be finally charged. In fact, based on Corollary 3.2, if the gross premium is always positive, we shall have the conditions

### v2

@@ 3.1 Equivalent martingale measures
Recall that $\{\mathrm{e}^{K(t)}\theta^{N_{t}}\mathrm{e}^{B(t)\lambda_{t}}~\mathrm{e}^{-\nu C_{t}}\mathrm{e}^{\phi\Lambda_{t}}\psi^{M_{t}}\}_{t\in{\cal T}}$ is a (local)-martingale. We suppose that some (suitable) square integrability conditions hold so that $\{\mathrm{e}^{K(t)}\theta^{N_{t}}\mathrm{e}^{B(t)\lambda_{t}}~\mathrm{e}^{-\nu C_{t}}\mathrm{e}^{\phi\Lambda_{t}}\psi^{M_{t}}\}_{t\in{\cal T}}$ is a square-integrable martingale. We now describe how the probability laws of $N_{t}$ and $C_{t}$ change after changing probability measures by the Esscher transform. Firstly, from Theorem 3.1, $\{\mathrm{e}^{K(t)}\theta^{N_{t}}\mathrm{e}^{B(t)\lambda_{t}}~\mathrm{e}^{-\nu C_{t}}\mathrm{e}^{\phi\Lambda_{t}}\psi^{M_{t}}\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P})$-martingale. Then, for all $t\in{\cal T}$,
@@
Consider the following (normalised) $\mathbb{F}$-adapted process $\{\zeta_{t}\}_{t\in{\cal T}}$:
@@
It is clear that by its definition in Eq. 3.9 and Theorem 3.1, $\{\zeta_{t}\}_{t\in{\cal T}}$ is an $(\mathbb{F},\mathbb{P})$-martingale. For each $t\in{\cal T}$, $\zeta_{t}>0$, $\mathbb{P}$-a.s. Furthermore, $\mathbb{E}_{0}[\zeta_{t}]=\zeta_{0}=1$. Consequently, a new probability measure $\tilde{\mathbb{P}}$, which is equivalent to the original probability measure $\mathbb{P}$ on ${\cal F}_{T}$, can be defined by putting:
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}\bigg|_{{\cal F}_{T}}:=\zeta_{T}.$
Given that $\{\zeta_{t}\}_{t\in{\cal T}}$ is a square-integrable $(\mathbb{F},\mathbb{P})$–martingale, then
$\mathbb{E}\left[\bigg|\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}\bigg|^{2}\right]=\mathbb{E}[|\zeta_{T}|^{2}]<\infty.$
Consequently, Definition 3.3 (ii) is satisfied.
We refer to related works employing the exponential change of measure: Palmowski and Rolski (2002) for Markov processes, Kallsen and Shiryaev (2002) via the cumulant process and Esscher transform, and Palmowski et al. (2025) for ruin probabilities with Hawkes arrivals.
3.2 Infinitesimal generator under the equivalent martingale measure
Let $\tilde{\mathcal{A}}$ denote the infinitesimal generator of the joint process $\{(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t)\}_{t\in{\cal T}}$ under the new probability measure $\tilde{\mathbb{P}}$. The key to study how the probability laws of $N_{t}$ and $C_{t}$ change after changing probability measures by the Esscher transform is to obtain the infinitesimal generator $\tilde{\mathcal{A}}$. This is to be achieved in Theorem 3.3 and Corollary 3.1 to be presented in the sequel. To do so we start with a technical lemma.
Lemma 3.1.
Assume that $\tilde{f}(\lambda,n,c,m,\Lambda,t)=\tilde{f}(\lambda,t)$ for all $n$, $c$, $m$ and $\Lambda$, and that $\mathrm{e}^{-B(t)\lambda_{t}}$ is a martingale. Consider $B(t)>0$ for some $t>0$. Then
$\tilde{\mathcal{A}}\tilde{f}(\lambda,0)=\frac{\mathcal{A}\bigl(\tilde{f}(\lambda,0)\,\mathrm{e}^{{-}B(t)\lambda_{t}}\bigr)}{\mathrm{e}^{-B(t)\lambda_{t}}}.$ (3.10)
Proof.
The generator of the process $(\lambda_{t},t)$ acting on a function $\tilde{f}(\lambda,t)$ with respect to the equivalent martingale probability measure is
$\tilde{\mathcal{A}}\tilde{f}(\lambda,0)=\lim_{t\downarrow 0}\frac{\tilde{\mathbb{E}}\!\left[\tilde{f}(\lambda_{t},t)\mid\lambda_{0}=\lambda\right]-\tilde{f}(\lambda,0)}{t}.$ (3.11)
We will use
$\frac{\mathrm{e}^{-B(t)\lambda_{t}}}{\mathbb{E}\!\left(\mathrm{e}^{-B(t)\lambda_{t}}\right)}$
as the Radon–Nikodym derivative to define the equivalent martingale probability measure, where $\mathbb{E}\!\left(\mathrm{e}^{-B(t)\lambda_{t}}\right)<\infty$. Hence, the expected value of $\tilde{f}(\lambda_{t},t)$ given $\lambda$ with respect to the equivalent martingale probability measure is
$\tilde{\mathbb{E}}\!\left[\tilde{f}(\lambda_{t},t)\mid\lambda_{0}=\lambda\right]=\frac{\mathbb{E}\!\left[\tilde{f}(\lambda_{t},t)\,\mathrm{e}^{-B(t)\lambda_{t}}\mid\lambda_{0}=\lambda\right]}{\mathbb{E}\!\left(\mathrm{e}^{-B(t)\lambda_{t}}\mid\lambda_{0}=\lambda\right)}.$ (3.12)
Since the denominator in (3.12) is a martingale, it becomes
$\tilde{\mathbb{E}}\!\left[\tilde{f}(\lambda_{t},t)\mid\lambda_{0}=\lambda\right]=\frac{\tilde{f}(\lambda,0)\,\mathrm{e}^{-B(t)\lambda}+\int_{0}^{t}\mathbb{E}\!\left[\mathcal{A}\tilde{f}(\lambda_{s},s)\,\mathrm{e}^{-B(t)\lambda_{s}}\mid\lambda_{0}=\lambda\right]\mathop{}\!\mathrm{d}s}{\mathrm{e}^{-B(t)\lambda}}.$ (3.13)
Substituting (3.13) into (3.11) gives
$\tilde{\mathcal{A}}\tilde{f}(\lambda,0)=\frac{1}{\mathrm{e}^{-B(t)\lambda}}\lim_{t\downarrow 0}\frac{1}{t}\int_{0}^{t}\mathbb{E}\!\left[\mathcal{A}\tilde{f}(\lambda_{s},s)\,\mathrm{e}^{-B(t)\lambda_{s}}\mid\lambda_{0}=\lambda\right]\mathop{}\!\mathrm{d}s.$ (3.14)
Therefore, by Dynkin’s formula (see Øksendal, 2013), (3.10) follows immediately. ∎
@@ 3.2 Infinitesimal generator under the equivalent martingale measure
[…]
$\hskip-14.22636pt\tilde{\mathcal{A}}\tilde{f}(\lambda,n,c,m,\Lambda,t)$
@@
$\quad+\theta\hat{j}(\nu)\hat{g}(-B(t))~\lambda\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}(\lambda+y,n+1,c+\xi,m,\Lambda,t)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right$]
$\quad+\tilde{\rho}(t)\left[\int_{0}^{\infty}\tilde{f}(\lambda+x,n,c,m+1,\Lambda,t)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right$]. (3.15)
@@
$\mathrm{e}^{K(t)}\theta^{n}\mathrm{e}^{B(t)\lambda}~\mathrm{e}^{-\nu c}\mathrm{e}^{\phi\Lambda}\psi^{m}\tilde{f}(\lambda,n,c,m,\Lambda,t),$
into the generator in Section 2.2.1 and from Lemma 3.1, we have
$\hskip-2.84544pt\tilde{\mathcal{A}}\tilde{f}(\lambda,n,c,m,\Lambda,t)$
$=\frac{\mathcal{A}\bigg\{\mathrm{e}^{K(t)}\theta^{n}\mathrm{e}^{B(t)\lambda}~\mathrm{e}^{-\nu c}\mathrm{e}^{\phi\Lambda}\psi^{m}\tilde{f}(\lambda,n,c,m,\Lambda,t)\bigg\}}{\mathrm{e}^{K(t)}\theta^{n}\mathrm{e}^{B(t)\lambda}~\mathrm{e}^{-\nu c}\mathrm{e}^{\phi\Lambda}\psi^{m}}$
[…]
$\quad+\lambda\left[\theta\hat{j}(\nu)\hat{g}(-B(t))\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}(\lambda+y,n+1,c+\xi,m,\Lambda,t)\frac{\mathrm{e}^{B(t)y}}{\hat{g}(-B(t))}\mathop{}\!\mathrm{d}G(y)\frac{\mathrm{e}^{-\nu\xi}}{\hat{j}(\nu)}\mathop{}\!\mathrm{d}J(\xi)-\tilde{f}\right$]
$\quad+\rho\left[\psi\hat{h}(-B(t))\int_{0}^{\infty}\tilde{f}(\lambda+x,n,c,m+1,\Lambda,t)\frac{\mathrm{e}^{B(t)x}}{\hat{h}(-B(t))}\mathop{}\!\mathrm{d}H(x)-\tilde{f}\right$].
@@
$\hskip-2.84544pt\tilde{\mathcal{A}}\tilde{f}(\lambda,n,c,m,\Lambda,t)$
[…]
$\quad+\lambda\left[\theta\hat{j}(\nu)\hat{g}(-B(t))\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}(\lambda+y,n+1,c+\xi,m,\Lambda,t)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right$]
$\quad+\rho\left[\psi\hat{h}(-B(t))\int_{0}^{\infty}\tilde{f}(\lambda+x,n,c,m+1,\Lambda,t)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right$]
@@
$\quad+\lambda\left[\theta\hat{j}(\nu)\hat{g}(-B(t))\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}(\lambda+y,n+1,c+\xi,m,\Lambda,t)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right$]
$\quad+\rho\left[\psi\hat{h}(-B(t))\int_{0}^{\infty}\tilde{f}(\lambda+x,n,c,m+1,\Lambda,t)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right$]
@@
$\quad+\theta\hat{j}(\nu)\hat{g}(-B(t))\lambda\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}(\lambda+y,n+1,c+\xi,m,\Lambda,t)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right$]
$\quad+\psi\hat{h}(-B(t))~\rho\left[\int_{0}^{\infty}\tilde{f}(\lambda+x,n,c,m+1,\Lambda,t)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right$].
@@
    The intensity process $\lambda_{t}$ has changed to $\tilde{\lambda}_{t}:=\theta\hat{j}(\nu)\hat{g}(-B(t))\,\lambda_{t}$, which is time-dependent;
@@
$\hskip-14.22636pt\tilde{\mathcal{A}}\tilde{f}(\tilde{\lambda},n,c,m,\tilde{\Lambda},t)$
@@
$\quad+\tilde{\lambda}\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}(\tilde{\lambda}+u,n+1,c+\xi,m,\tilde{\Lambda},t)\mathop{}\!\mathrm{d}\tilde{G}\left(\frac{u}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t\right)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right$]
$\quad+\tilde{\rho}(t)\left[\int_{0}^{\infty}\tilde{f}(\tilde{\lambda}+v,n,c,m+1,\tilde{\Lambda},t)\mathop{}\!\mathrm{d}\tilde{H}\left(\frac{v}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t\right)-\tilde{f}\right$]. (3.16)
@@
$\hskip-14.22636pt\tilde{\mathcal{A}}\tilde{f}(\tilde{\lambda},n,c,m,\tilde{\Lambda},t)$
@@
$\quad+\tilde{\lambda}\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}(\tilde{\lambda}+\theta\hat{j}(\nu)\hat{g}(-B(t))y,n+1,c+\xi,m,\tilde{\Lambda},t)\mathop{}\!\mathrm{d}\tilde{G}(y;t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)-\tilde{f}\right$]
$\quad+\tilde{\rho}(t)\left[\int_{0}^{\infty}\tilde{f}(\tilde{\lambda}+\theta\hat{j}(\nu)\hat{g}(-B(t))x,n,c,m+1,\tilde{\Lambda},t)\mathop{}\!\mathrm{d}\tilde{H}(x;t)-\tilde{f}\right$].
@@
$a\$$\rightarrow\ \theta\hat{j}(\nu)\hat{g}(-B(t))~a;$
$\rho\$$\rightarrow\ \psi\hat{h}(-B(t))~\rho;$
$h(v)\$$\rightarrow\ \frac{\tilde{h}\left(\frac{v}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t\right)}{\theta\hat{j}(\nu)\hat{g}(-B(t))};$
$g(u)\$$\rightarrow\ \frac{\tilde{g}\left(\frac{u}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t\right)}{\theta\hat{j}(\nu)\hat{g}(-B(t))};$
$j(\xi)\$$\rightarrow\ \tilde{j}(\xi),$
@@
$\hskip-14.22636pt\tilde{\mathcal{A}}\tilde{f}(\tilde{\lambda},n,c,m,\tilde{\Lambda},t)$
@@
$\quad+\tilde{\lambda}\left[\int_{0}^{\infty}\int_{0}^{\infty}\tilde{f}(\tilde{\lambda}+u,n+1,c+\xi,m,\tilde{\Lambda},t)\mathop{}\!\mathrm{d}\tilde{J}(\xi)~\frac{\tilde{g}(\frac{u}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t)}{\theta\hat{j}(\nu)\hat{g}(-B(t))}\mathop{}\!\mathrm{d}u-\tilde{f}\right$]
$\quad+\tilde{\rho}(t)\left[\int_{0}^{\infty}\tilde{f}(\tilde{\lambda}+v,n,c,m+1,\tilde{\Lambda},t)\frac{\tilde{h}(\frac{v}{\theta\hat{j}(\nu)\hat{g}(-B(t))};t)}{\theta\hat{j}(\nu)\hat{g}(-B(t))}\mathop{}\!\mathrm{d}v-\tilde{f}\right$].
@@
Note that the parameters of a CDCP under the original measure $\mathbb{P}$ are all constants, whereas the parameters of a CDCP under the new measure $\tilde{\mathbb{P}}$ become time-varying (except for the intensity-decay rate $\delta$), i.e., a time-inhomogeneous CDCP.
@@
In a hardening reinsurance market, it may be reasonable for a reinsurer to calculate the premium of a catastrophe reinsurance contract in a more conservative and prudent way. To illustrate this situation in our current modelling set up, certain assumptions are imposed on $\theta,\psi,\nu,B(t)$ such that there would be more external-exciting jump arrivals in a given period of time, a higher value of self-exciting intensity, higher values of their jump sizes, as well as a larger size of claim or loss. To take on these additional risks in the hardening market, a reinsurer requires a higher compensation. Furthermore, it may be reasonable to postulate that the reinsurer aims to maximise their shareholders’ wealth via earning profits. To this end, the reinsurer may charge a higher premium than the “break-even” premium. The latter is given by the expected claims evaluated under the original (real-world) probability measure. In the hardening market, $\theta,\psi,\nu,B(t)$ may be thought of as security loading factors to be evaluated by which positive gross premium should be finally charged. In fact, based on Corollary 3.2, if the gross premium is always positive, we shall have the conditions